OpenStudy (anonymous):

Find the sum of first 35 terms of the series whose pth term is p/7 +2

OpenStudy (vishweshshrimali5):

Are you sure it is p ?

OpenStudy (vishweshshrimali5):

Ohh okay

OpenStudy (vishweshshrimali5):

\[\large{a_p = \cfrac{p}{7} + 2}\]

OpenStudy (anonymous):

a+(p-1)d =.........

OpenStudy (vishweshshrimali5):

First you have to be sure that it is an AP

OpenStudy (vishweshshrimali5):

To check that calculate the difference between two consecutive terms

OpenStudy (anonymous):

It is an AP , it is given under AP, but let us check anyway

OpenStudy (vishweshshrimali5):

\[\large{a_p - a_{p-1} = \cfrac{p}{7}-\cfrac{p-1}{7}}\] \[\large{\implies a_p - a_{p-1} = \cfrac{p - p +1}{7} = \cfrac{1}{7}}\] Now this is independent of p, thus it is an ap

OpenStudy (anonymous):

why p/7 and p-1/7

OpenStudy (anonymous):

i mean how

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @No.name the series whose pth term is p/7 +2 \(\color{blue}{\text{End of Quote}}\) If pth term was p/7 + 2, then (p-1)th term would have been (p-1)/7 + 2 I cancelled out the 2 as both a_p and a_(p-1) had "+2"

OpenStudy (anonymous):

uhm yess

OpenStudy (vishweshshrimali5):

Good. Now, if this is an ap, then the difference of the consecutive terms = common difference = d

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

I got it here, the sollution http://yourmathguru.com/forums/topic/summation-4

OpenStudy (anonymous):

the last sollution

OpenStudy (vishweshshrimali5):

That was the same as the one using \(\large{\sum}\)

OpenStudy (anonymous):

Oh i have to go , will contuinue tommorow bye!

OpenStudy (vishweshshrimali5):

bye